# A New Algorithm for Zero-Sum Linear-Quadratic Stochastic Differential Games in Infinite Horizons

### v1

@@ Abstract
We propose a new algorithm for Zero-Sum Linear-Quadratic Stochastic Differential Games by dissecting their inherent structures. Specifically, we construct dual-layer iterative matrix-increasing sequences, which reformulate the original problem into a set of mutually interconnected subproblems. By sequentially computing the stabilizing solutions to the associated algebraic Riccati equations within each subproblem, we derive the stabilizing solutions for the original problem and rigorously establish the convergence of the proposed algorithm. Numerical simulations further validate the effectiveness of the method. This work extends classical setting and provides the first complete unified numerical framework for solving a broader class of stochastic Game-Theoretic algebraic Riccati equations.
Keywords: Iteration Algorithm; Zero-Sum Linear-Quadratic Stochastic Differential Games; Game-Theoretic algebraic Riccati equations; Stabilizing Solutions
MSCcodes: 65F45, 91A15, 93D15
@@ 1 Introduction
Zero-Sum Stochastic Differential games provide a rigorous mathematical framework for modeling adversarial decision-making in stochastic dynamic systems. Building on the foundational work of [17] and [24] in differential games, [15] extended the deterministic framework developed by [20] to stochastic settings. Within this framework, the Zero-Sum Linear-Quadratic Stochastic Differential Game (ZSLQSDG) serves as a pivotal simplified model. It circumvents the significant challenges associated with solving the highly nonlinear Hamilton-Jacobi-Bellman-Isaacs equation, making it a key vehicle for both theoretical exploration and applied research. To date, significant advancements have been achieved in the theoretical study of ZSLQSDG; see, for example, [23, 27, 32, 31, 26, 34, 33, 29].
@@
This paper focuses on the classic infinite-horizon ZSLQSDG with constant coefficients. For infinite-horizon problems–whether in deterministic or stochastic cases–only the stabilizing solution is of practical interest among all possible solutions. Specifically for this problem, [31] proves that the existence of closed-loop saddle points is characterized by the solvability of an algebraic Riccati equation (AREs) with an indefinite quadratic term under a certain stabilizing condition. In other words, solving a ZSLQSDG problem is equivalent to solving its associated AREs.
@@
In the existing literature, AREs featuring indefinite quadratic terms are commonly referred to as Game-Theoretic algebraic Riccati equations (GTAREs). For further relevant studies, see [6, 25, 10, 3, 2, 12, 5, 33] and their references. Due to the indefiniteness of their quadratic terms, the traditional numerical algorithms for solving conventional AREs are no longer applicable. For deterministic GTAREs, [21] proposed a recursive method for the positive stabilizing solution. Unlike deterministic systems, stochastic systems with state-dependent noise complicate stabilizing solution computation. This complexity renders traditional deterministic methods inapplicable to the above cases. To address this, [14] developed an algorithm for GTAREs arising from a specific LQZSDG with state-dependent noise. This approach was later extended by [11] to stochastic $H_{\infty}$ problems involving state- and control-dependent noise, and further by [7] to stochastic $H_{\infty}$ problems involving state-, control-, and disturbance-dependent noise. [19] investigated a class GTAREs arising in stochastic $H_{\infty}$ problems with state-dependent noise. The study explored three effective approaches for computing stabilizing solutions of these equations, and further performed numerical comparisons and analyses of the iterative methods among them. Further computational studies can be found in [4, 9, 18, 8, 1, 30] and their references.
To date, no iterative algorithm has been reported in the existing literature for classical ZSLQSDG with state- and both-player-dependent noise. This paper develops a new algorithm for ZSLQSDG by dissecting their inherent structures. we construct dual-layer matrix iteration sequences via defect correction method proposed in [22], which reformulate the ZSLQSDG problem into a series of interconnected bilevel subproblems. These subproblems are solved sequentially by identifying the maximal solutions to their associated AREs, ultimately yielding stabilizing solutions for the original problem. Importantly, this study first extends the classical framework and develops a unified algorithm applicable to three classes of problems: deterministic GTAREs, stochastic $H_{\infty}$-type GTAREs, and the more general stochastic GTAREs associated with the classical ZSLQSDG.
@@
Remaining Content Outline: Section 2 covers the mathematical framework of classical ZSLQSDG, a decoupled representation method, and the algorithm iteration process. Section 3 presents the main results, including the construction logic of the inner-outer matrix iteration sequence, the proofs of existence, uniqueness, and boundedness of the stabilizing solutions within each subproblem, as well as the analysis of algorithm convergence. Section 4 details the content related to numerical experiments.
@@ 2.1 Notation
Let us first introduce the following notation that will be used below:
  • $\mathbb{R}^{n}$: $n$-dimensional Euclidean space; $\mathbb{C}^{-}$: complex numbers with negative real part; $\mathbb{R}^{n\times m}$: the set of all $n\times m$ real matrices; $\mathbb{S}^{n}$: the set of all $n\times n$ symmetric matrices; $\overline{\mathbb{S}}_{+}^{n}$: the set of all $n\times n$ symmetric positive semi-definite matrices; $\mathbb{S}_{+}^{n}$: the set of all $n\times n$ symmetric positive definite matrices.
  • $I_{n}$: the identity matrix of size $n$; $\mathbb{O}_{n\times m}$: the null matrix of size $n\times m$. It can be simplified as 0 when no ambiguity is generated; $\mathbb{O}_{m\times n}^{\times(l)}$ denotes the object represented by $\mathbb{O}_{m\times n}$ arranged repeatedly for $l$ times.
  • $A^{\top}$: the transpose of the matrix $A$; $A^{\dagger}$: the Moore-Penrose pseudoinverse of the matrix $A$; $\mathrm{Tr}(\cdot)$: the trace of a square matrix; $\langle\cdot,\cdot\rangle$: the inner product on a Hilbert space. In particular, the usual inner product on $\mathbb{R}^{n\times m}$ is given by $\langle A,B\rangle\mapsto\mathrm{Tr}(A^{\top}B)$;
  • If $A\in\mathbb{S}_{+}^{n}$ (resp., $A\in\overline{\mathbb{S}}_{+}^{n}$) is a symmetric positive definite (resp., symmetric positive semi-definite) matrix, we write $A\succ 0$ (resp., $A\succeq 0$). For any $A,B\in\mathbb{S}^{n}$, we use the notation $A\succ B$ (resp., $A\succeq B$) to indicate that $A-B\succ 0$ (resp., $A-B\succeq 0$).
@@
$L_{\mathbb{F}}^{2}(\mathbb{H})=\left\{\varphi:[0,\infty)\times\Omega\rightarrow\mathbb{H}\mid\varphi\in\mathbb{F}\text{ and }\mathbb{E}\int_{0}^{\infty}|\varphi(t)|^{2}\,dt<\infty\right\},\quad\mathcal{U}_{i}=L_{\mathbb{F}}^{2}(\mathbb{R}^{m_{i}}),\,i=1,2.$
$\mathcal{X}_{\mathrm{loc}}[0,\infty)=\left\{\varphi:[0,\infty)\times\Omega\rightarrow\mathbb{R}^{n}\mid\varphi\in\mathbb{F}\text{ is continuous, and }\mathbb{E}\left[\sup_{0\leq t\leq T}|\varphi(t)|^{2}\right]<\infty\text{ for every }T>0\right\},$
$\mathcal{X}[0,\infty)=\left\{\varphi\in\mathcal{X}_{\mathrm{loc}}[0,\infty)\mid\mathbb{E}\int_{0}^{\infty}|\varphi(t)|^{2}\,dt<\infty\right\}.$
@@ 2.2 Zero-Sum Linear-Quadratic Stochastic Differential Games in Infinite Horizons
Let $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ be a complete filtered probability space on which a r-dimensional standard Brownian motion $W=\{w_{1}(t),\dots,w_{r}(t);t\geq 0\}$ is defined with $\mathbb{F}=\{\mathcal{F}_{t}\}_{t\geq 0}$ being the usual augmentation of the natural filtration generated by $W$. We consider the following controlled linear stochastic differential equation (SDE) on $0,\infty)$:
$\begin{cases}dX(t)=[AX(t)+B_{1}u_{1}(t)+B_{2}u_{2}(t)]dt+\sum_{l=1}^{r}[C_{l}X(t)+D_{l,1}u_{1}(t)+D_{l,2}u_{2}(t)]dw_{l}(t)\\ X(0)=x\end{cases},$ (1)
@@
$J(x;u_{1},u_{2})\triangleq\mathbb{E}\int_{0}^{\infty}\Bigg[\Bigg\langle\begin{pmatrix}Q&S_{1}^{\top}&S_{2}^{\top}\\ S_{1}&R_{11}&R_{12}\\ S_{2}&R_{21}&R_{22}\end{pmatrix}\begin{pmatrix}X(t)\\ u_{1}(t)\\ u_{2}(t)\end{pmatrix},\begin{pmatrix}X(t)\\ u_{1}(t)\\ u_{2}(t)\end{pmatrix}\Bigg\rangle\Bigg]dt,$ (2)
@@
For $(x,u_{1},u_{2})\in\mathbb{R}^{n}\times\mathcal{U}_{1}\times\mathcal{U}_{2}$, the solution $X(\cdot;x,u_{1},u_{2})$ to the SDE in (1) may only exist in $\mathcal{X}_{loc}[0,\infty)$, which renders $J(x;u_{1},u_{2})$ ill-defined. We define the set of admissible controls as $\mathcal{U}_{ad}(x)\triangleq\{(u_{1},u_{2})\in\mathcal{U}_{1}\times\mathcal{U}_{2}\mid X(\cdot;x,u_{1},u_{2})\in\mathcal{X}[0,\infty)\}$. A pair $(u_{1},u_{2})\in\mathcal{U}_{ad}(x)$ is called an admissible control pair for the initial state $x$, and the corresponding $X(\cdot;x,u_{1},u_{2})$ is referred to as the admissible state process. In this case, $J(x,u_{1},u_{2})$ is clearly well-defined.
In this zero-sum game, Player 1 (the maximizer) selects control $u_{1}$ to maximize (2), while Player 2 (the minimizer) chooses $u_{2}$ to minimize the same function. The problem is to find an admissible control pair $(u_{1}^{*},u_{2}^{*})$ that both players can accept. We denote the above-mentioned problem as $({SDG})^{\,0}_{\infty}$ for short. For a description of the $({SDG})^{\,0}_{\infty}$ problem, refer to [28] . More detailed information can be found therein.
The $({SDG})^{\,0}_{\infty}$ problem corresponds to following stochastic GTAREs:
$\begin{cases}Q(P)-S(P)^{\top}R(P)^{\dagger}S(P)=0\\ \mathcal{R}(S(P))\subseteq\mathcal{R}(R(P))\\ R_{11}(P)\succeq 0,\quad R_{22}(P)\preceq 0\end{cases},$ (3)
where $\mathcal{R}(M)$ denotes the range of a matrix or an operator $M$,
$Q(P)=PA+A^{\top}P+\sum_{l=1}^{r}C_{l}^{\top}PC_{l}+Q,$
@@
For the stochastic GTAREs (3), the general theoretical framework of $({SDG})^{\,0}_{\infty}$ only requires that $R(P)$ satisfies $\mathcal{R}(S(P))\subseteq\mathcal{R}(R(P))$ in [28]. However, in our subsequent discussion, we impose the slightly stronger condition that $R(P)$ be of full rank. This requirement is well-motivated as it allows us to focus on the core structural properties of the ZSLQSDG. We therefore restrict our analysis to the case where $R(P)$ is full rank.
@@
The system (1) donated as $(A,C_{1},\dotsb,C_{r})$ is called mean-square stable such that the solution of the following SDE:
$\begin{cases}dX(t)=AX(t)dt+\sum_{l=1}^{r}C_{l}X(t)dw_{l}(t)\\ X(0)=x\end{cases},$ (4)
satisfies $\lim_{t\to\infty}\mathbb{E}[X(t)^{\top}X(t)]=0$ for every initial state $x\in\mathbb{R}^{n}$.
A concept related to the above admissible control pair is that of mean-square stabilizers. We define the set of all such stabilizers associated with the system (1) as $\mathcal{K}=$
$\left\{(\varTheta_{1},\varTheta_{2})\in\mathbb{R}^{m_{1}\times n}\times\mathbb{R}^{m_{2}\times n}\ \middle|\ \begin{aligned} &\text{The system}\\ &(A+B_{1}\varTheta_{1}+B_{2}\varTheta_{2},C_{1}+D_{1,1}\varTheta_{1}+D_{1,2}\varTheta_{2},\dotsb,C_{r}+D_{r,1}\varTheta_{1}+D_{r,2}\varTheta_{2})\\ &\text{ is mean-square stable}.\end{aligned}\right\},$
Let $(\varTheta_{1},\varTheta_{2})\in\mathcal{K}$, the feedback control $U(\cdot)=\begin{bmatrix}\varTheta_{1}X(\cdot)\\ \varTheta_{2}X(\cdot)\end{bmatrix}$ is called mean-square stabilizing. We refer to [25, 28] for the above definition.
Definition 2.3 ([*]Sun2020_book).
@@
Theorem 2.6.7 in [28] is a direct conclusion pertaining to the relationship between $({SDG})^{\,0}_{\infty}$ and stochastic GTAREs. Leveraging this theoretical result, we convert $({SDG})^{\,0}_{\infty}$ into an equivalent problem of finding the stabilizing solution $P$ to the stochastic GTARE (3).
@@
$\begin{cases}dX(t)=[A_{L}X(t)+B_{1}v_{1}(t)]dt+\sum_{l=1}^{r}[C_{lL}X(t)+D_{l,1}v_{1}(t)]dw_{l}(t)\\ x(0)=x\end{cases}$ (5)
@@
$J_{L}(x;v_{1})\triangleq\mathbb{E}\int_{0}^{\infty}\Bigg[\Bigg\langle\begin{pmatrix}Q_{L}&S^{\top}_{L}\\ S_{L}&R_{11}\end{pmatrix}\begin{pmatrix}X(t)\\ v_{1}(t)\end{pmatrix},\begin{pmatrix}X(t)\\ v_{1}(t)\end{pmatrix}\Bigg\rangle\Bigg]dt,$ (6)
@@
$\begin{cases}A_{L}=A+B_{1}K_{1}(0)+B_{2}K_{2}(0)+B_{2}L,\,C_{lL}=C_{l}+D_{l,1}K_{1}(0)+D_{l,2}K_{2}(0)+D_{l,2}L,&1\leq l\leq r\\ Q_{L}=Q-S^{\top}(0)R(0)^{-1}S(0)+L^{\top}R_{22}L,\quad S_{L}=R_{12}L\end{cases}.$
@@
$PA_{L}+A_{L}^{\top}P+\sum_{l=1}^{r}C_{lL}^{\top}PC_{lL}+Q_{L}$ (7)
$-\left(B_{1}^{\top}P+\sum_{l=1}^{r}D_{l,1}^{\top}PC_{lL}+S_{L}\right)^{\top}\left(R_{11}+\sum_{l=1}^{r}D_{l,1}^{\top}PD_{l,1}\right)^{-1}\left(B_{1}^{\top}P+\sum_{l=1}^{r}D_{l,1}^{\top}PC_{lL}+S_{L}\right)=0.$

### v2

# A Unified Computational Approach for Zero-Sum Linear-Quadratic Stochastic Differential Games in Infinite Horizons

@@ Abstract
This paper proposes a new method for finding closed-loop saddle points in zero-sum linear-quadratic stochastic differential games by decoupling their inherent structure. Specifically, we develop a nested iterative scheme that constructs a monotonically increasing sequence of matrices, thereby decomposing the original problem into interconnected subproblems. By sequentially computing the stabilizing solutions to the algebraic Riccati equations within each subproblem, we obtain the stabilizing solution to the original problem and rigorously establish the convergence of the iterative sequence. A numerical example further validates the effectiveness of the proposed method. To the best of our knowledge, this work extends the classical setting and provides the first general-purpose computational approach for this class of problems.
Keywords: Computational Approach, Stabilizing Solution, Zero-Sum Linear-Quadratic Stochastic Differential Games, Game-Theoretic Algebraic Riccati Equations
@@ 1 Introduction
Zero-sum stochastic differential games provide a rigorous mathematical framework for modeling adversarial decision-making in stochastic dynamic systems. Building on the foundational work of [Isaacs1965] and [Pontryagin1967], [Fleming1989] extended the deterministic framework of [Evans1984] to stochastic settings. Within this framework, the zero-sum linear-quadratic stochastic differential games (ZSLQSDG) serve as a canonical simplified model. This model circumvents the significant challenges associated with solving the highly nonlinear Hamilton-Jacobi-Bellman-Isaacs equation, making it a key vehicle for both theoretical and applied research. To date, significant advancements have been achieved in the theoretical study of ZSLQSDG; see, for example, [Mou2006, Sun2014, Yu2015, Sun2016, Sun2021, Zhou2022, Zhang2024, Sun2024].
@@
This paper focuses on the infinite-horizon ZSLQSDG with constant coefficients. For infinite-horizon problems, whether in deterministic or stochastic cases, only the stabilizing solution is of practical interest among all possible solutions. Specifically for this problem, [Sun2016] proves that the existence of closed-loop saddle points is characterized by the solvability of an algebraic Riccati equation (ARE) with a sign-indefinite quadratic term under a certain stabilizing condition. In other words, finding the closed-loop saddle point of the problem requires computing the stabilizing solution to its associated ARE.
@@
ARE featuring a sign-indefinite quadratic term is commonly referred to as Game-Theoretic algebraic Riccati equation (GTARE). For further relevant studies, see [Dragan1997, Rami2000, Dragan2001, Damm2004book, Abou2012book, Dragan2013book, Dragan2020, Zhang2024] and their references. Due to the indefiniteness of its quadratic term, the traditional numerical algorithms for solving conventional ARE are no longer applicable. For deterministic GTARE, [Lanzon2008] proposed a recursive method for the positive stabilizing solution. Unlike deterministic systems, stochastic systems with state-dependent noise complicate stabilizing solution computation. To address this, [Feng2010] developed a method for GTARE arising from a specific ZSLQSDG with state-dependent noise. This approach was later extended by [Dragan2011] to stochastic $H_{\infty}$ problems involving state-dependent and control-dependent noise, and further by [Dragan201501] to stochastic $H_{\infty}$ problems involving state-dependent, control-dependent and disturbance-dependent noise. [Ivanov2018] investigated a class of GTAREs arising in stochastic $H_{\infty}$ problems with state-dependent noise. The study explored three effective approaches for computing the stabilizing solutions of these equations and performed numerical comparisons of the iterative methods. Further relevant computational studies can be found in [Dragan2008, Dragan201502, Ivanov2015, Dragan2017, Aberkane2023, Sun2023] and their references.
To the best of our knowledge, no computational method has been reported in the literature for ZSLQSDG with state-dependent and both-control-dependent noise. This paper proposes a new method for ZSLQSDG by decoupling its inherent structure. Specifically, we construct a monotonically increasing sequence of matrices by extending the defect correction method proposed in [Mehrmann1988], thereby decomposing the ZSLQSDG problem into a series of interconnected subproblems. These subproblems are solved sequentially by identifying the stabilizing solutions to their associated AREs, ultimately yielding the stabilizing solutions for the original problem. The main contributions of this work are summarized as follows:
  • First General Computational Approach: Previous research has only tackled special cases of this problem, this work overcomes the inherent coupling challenges and provides the first general computational method for finding the closed-loop saddle points of ZSLQSDG.
  • Extended Applicability: The computational method extends the classical setting and is applicable to three categories of GTAREs: deterministic GTAREs, stochastic $H_{\infty}$-type GTAREs, and more general stochastic GTAREs arising from ZSLQSDG.
  • Theoretical Guarantees: This study establishes rigorous theoretical guarantees for the nested iterative scheme, including the existence, uniqueness and boundedness of the stabilizing solutions to AREs with sign-definite quadratic terms, along with a comprehensive convergence analysis.
@@
The remainder of this paper is organized as follows. Section 2 covers the mathematical framework of ZSLQSDG, a decoupled representation method, and the iteration process of the computational method. Section 3 presents the main results, including the key structural properties of the problem, the criteria for the existence and uniqueness of the stabilizing solutions to the subproblems, and the proofs for the boundedness and convergence of the iterative sequence. Section 4 details a numerical example.
@@ 2.1 Notation
Let us first introduce the following notation:
  • $\mathbb{R}^{n}$: $n$-dimensional real Euclidean space; $\mathbb{C}^{-}$: the set of complex numbers with negative real part; $\mathbb{R}^{n\times m}$: the space of $n\times m$ real matrices; $\mathbb{S}^{n}$: the set of all $n\times n$ symmetric matrices; $\overline{\mathbb{S}}_{+}^{n}$: the set of all $n\times n$ symmetric positive semi-definite matrices; $\mathbb{S}_{+}^{n}$: the set of all $n\times n$ symmetric positive definite matrices.
  • $I_{n}$: the identity matrix of size $n$; $\mathbb{O}_{n\times m}$: the null matrix of size $n\times m$. It can be simplified as 0 when no ambiguity is generated; $A_{m\times n}^{\times(l)}$ denotes the object represented by $A_{m\times n}$ arranged repeatedly for $l$ times.
  • $A^{\top}$: the transpose of the matrix $A$; $A^{\dagger}$: the Moore-Penrose pseudoinverse of the matrix $A$; $\langle\cdot,\cdot\rangle$: the inner product on a Hilbert space. If $A\in\mathbb{S}_{+}^{n}$ (resp., $A\in\overline{\mathbb{S}}_{+}^{n}$), we write $A\succ 0$ (resp., $A\succeq 0$). For any $A,B\in\mathbb{S}^{n}$, we use the notation $A\succ B$ (resp., $A\succeq B$) to indicate that $A-B\succ 0$ (resp., $A-B\succeq 0$).
@@ 2.2 Zero-Sum Linear-Quadratic Stochastic Differential Games in Infinite Horizons
Let $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ be a complete filtered probability space on which a r-dimensional standard Brownian motion $W=\{W(t)^{\top}=(w_{1}(t),\dotsb,w_{r}(t));t\geq 0\}$ is defined with $\mathbb{F}=\{\mathcal{F}_{t}\}_{t\geq 0}$ being the usual augmentation of the natural filtration generated generated by $W$. We consider the following controlled linear stochastic differential equation (SDE) on $0,\infty)$:
$\begin{cases}&dX(t)=[AX(t)+B_{1}u_{1}(t)+B_{2}u_{2}(t)]dt\\ &\quad\quad\quad+\sum_{l=1}^{r}[C_{l}X(t)+D_{l,1}u_{1}(t)+D_{l,2}u_{2}(t)]dw_{l}(t)\\ &X(0)=x\end{cases},$ (1)
@@
$J(x;u_{1},u_{2})=\mathbb{E}\int_{0}^{\infty}\Bigg[\Bigg\langle\begin{pmatrix}Q&S_{1}^{\top}&S_{2}^{\top}\\ S_{1}&R_{11}&R_{12}\\ S_{2}&R_{21}&R_{22}\end{pmatrix}\begin{pmatrix}X(t)\\ u_{1}(t)\\ u_{2}(t)\end{pmatrix},\begin{pmatrix}X(t)\\ u_{1}(t)\\ u_{2}(t)\end{pmatrix}\Bigg\rangle\Bigg]dt,$ (2)
@@
For $(x,u_{1},u_{2})\in\mathbb{R}^{n}\times\mathcal{U}_{1}\times\mathcal{U}_{2}$, the solution $X(\cdot;x,u_{1},u_{2})$ to the SDE in (1) may only exist in $\mathcal{X}_{loc}[0,\infty)$, which renders $J(x;u_{1},u_{2})$ ill-defined. We define the set of admissible controls as $\mathcal{U}_{ad}(x)=\{(u_{1},u_{2})\in\mathcal{U}_{1}\times\mathcal{U}_{2}\mid X(\cdot;x,u_{1},u_{2})\in\mathcal{X}[0,\infty)\}$. A pair $(u_{1},u_{2})\in\mathcal{U}_{ad}(x)$ is called an admissible control pair for the initial state $x$, and the corresponding $X(\cdot;x,u_{1},u_{2})$ is referred to as the admissible state process. In this case, $J(x,u_{1},u_{2})$ is clearly well-defined.
In this zero-sum game, Player 1 (the maximizer) selects control $u_{1}$ to maximize (2), while Player 2 (the minimizer) chooses $u_{2}$ to minimize the same function. The problem is to find an admissible control pair $(u_{1}^{*},u_{2}^{*})$ that both players can accept. We denote the above-mentioned problem as $({SDG})^{\,0}_{\infty}$ for short. For a description of the $({SDG})^{\,0}_{\infty}$ problem, refer to [Sun2020_book] . More detailed information can be found therein.
The $({SDG})^{\,0}_{\infty}$ problem corresponds to following stochastic GTARE:
$\begin{cases}Q(P)-S(P)^{\top}R(P)^{\dagger}S(P)=0\\ \mathcal{R}(S(P))\subseteq\mathcal{R}(R(P))\\ R_{11}(P)\preceq 0,\quad R_{22}(P)\succeq 0\end{cases},$ (3)
where $\mathcal{R}(M)$ denotes the range of a matrix $M$,
$Q(P)=PA+A^{\top}P+\sum_{l=1}^{r}C_{l}^{\top}PC_{l}+Q,$ (4)
@@
For the stochastic GTARE (3), the general theoretical framework of $({SDG})^{\,0}_{\infty}$ only requires that $R(P)$ satisfies $\mathcal{R}(S(P))\subseteq\mathcal{R}(R(P))$ in [Sun2020_book]. However, in our subsequent discussion, we impose the slightly stronger condition that $R(P)$ be of full rank. This requirement is well-motivated as it allows us to focus on the core structural properties of the $({SDG})^{\,0}_{\infty}$. We therefore restrict our analysis to the case where $R(P)$ is full rank.
@@
The system
$\begin{cases}dX(t)=AX(t)dt+\sum_{l=1}^{r}C_{l}X(t)dw_{l}(t)\\ X(0)=x\end{cases},$
donated as $(A,C_{1},\dotsb,C_{r})$ is called mean-square stable such that the solution satisfies
$\lim_{t\to\infty}\mathbb{E}[X(t)^{\top}X(t)]=0\,\,\text{for every initial state}\,\,x\in\mathbb{R}^{n}.$
A concept related to the above admissible control pair is that of mean-square stabilizers. We define the set of all such stabilizers associated with the system (1) as
$\mathcal{K}=\left\{\begin{aligned} &\text{All of}\,\,\,(\varTheta_{1},\varTheta_{2})\in\mathbb{R}^{m_{1}\times n}\times\mathbb{R}^{m_{2}\times n}\,\text{such that the system}\,(A+B_{1}\varTheta_{1}+B_{2}\varTheta_{2},\\ &C_{1}+D_{1,1}\varTheta_{1}+D_{1,2}\varTheta_{2},\dotsb,C_{r}+D_{r,1}\varTheta_{1}+D_{r,2}\varTheta_{2})\text{ is mean-square stable}.\end{aligned}\right\}.$
Let $(\varTheta_{1},\varTheta_{2})\in\mathcal{K}$, the feedback control $U(\cdot)=\begin{bmatrix}\varTheta_{1}X(\cdot)\\ \varTheta_{2}X(\cdot)\end{bmatrix}$ is called mean-square stabilizing. Moreover, for $\varTheta_{i}\in\mathbb{R}^{m_{i}\times n}(i=1,2)$, we let
$\mathcal{K}_{1}(\varTheta_{2})=\left\{\varTheta_{1}\in\mathbb{R}^{m_{1}\times n}:(\varTheta_{1},\varTheta_{2})\in\mathcal{K}\right\},$
$\mathcal{K}_{2}(\varTheta_{1})=\left\{\varTheta_{2}\in\mathbb{R}^{m_{2}\times n}:(\varTheta_{1},\varTheta_{2})\in\mathcal{K}\right\}.$
We refer to [Rami2000, Sun2020_book] for the above definition.
Definition 2.3 ([Sun2020_book]).
A 2-tuple pair $(\varTheta_{1}^{*},\varTheta_{2}^{*})\in\mathbb{R}^{m_{1}\times n}\times\mathbb{R}^{m_{2}\times n}$ is called a closed-loop saddle point of Problem $({SDG})^{\,0}_{\infty}$ if $(\varTheta_{1}^{*},\varTheta_{2}^{*})\in\mathcal{K}$ and
$J(x,\varTheta_{1}X(\cdot),\varTheta_{2}^{*}X(\cdot))\leq J(x,\varTheta_{1}^{*}X(\cdot),\varTheta_{2}^{*}X(\cdot))\leq J(x,\varTheta_{1}^{*}X(\cdot),\varTheta_{2}X(\cdot)),$
for all $x\in\mathbb{R}^{n},(\varTheta_{1},\varTheta_{2})\in\mathcal{K}_{1}(\varTheta_{2}^{*})\times\mathcal{K}_{2}(\varTheta_{1}^{*})$.
Definition 2.4 ([Sun2020_book]).
@@
Theorem 2.6.7 in [Sun2020_book] directly establishes the relationship between $({SDG})^{\,0}_{\infty}$ and the stochastic GTARE. Based on this result, we transform the problem of finding the closed-loop saddle point of $({SDG})^{\,0}_{\infty}$ into the equivalent problem of finding the stabilizing solution $P$ to the stochastic GTARE (3).
@@
$\begin{cases}dX(t)=[A_{L}X(t)+B_{1}v_{1}(t)]dt+\sum_{l=1}^{r}[C_{lL}X(t)+D_{l,1}v_{1}(t)]dw_{l}(t)\\ x(0)=x\end{cases}$
@@
$J_{L}(x;v_{1})=\mathbb{E}\int_{0}^{\infty}\Bigg[\Bigg\langle\begin{pmatrix}Q_{L}&S^{\top}_{L}\\ S_{L}&R_{11}\end{pmatrix}\begin{pmatrix}X(t)\\ v_{1}(t)\end{pmatrix},\begin{pmatrix}X(t)\\ v_{1}(t)\end{pmatrix}\Bigg\rangle\Bigg]dt,$
@@
$\begin{cases}&A_{L}=A+B_{1}K_{1}(0)+B_{2}K_{2}(0)+B_{2}L\\ &C_{lL}=C_{l}+D_{l,1}K_{1}(0)+D_{l,2}K_{2}(0)+D_{l,2}L,1\leq l\leq r\\ &Q_{L}=Q-S^{\top}(0)R(0)^{-1}S(0)+L^{\top}R_{22}L,\quad S_{L}=R_{12}L\end{cases}.$
@@
$PA_{L}+A_{L}^{\top}P+\sum_{l=1}^{r}C_{lL}^{\top}PC_{lL}+Q_{L}-(B_{1}^{\top}P+\sum_{l=1}^{r}D_{l,1}^{\top}PC_{lL}+S_{L})^{\top}$ (5)
$\quad\times(R_{11}+\sum_{l=1}^{r}D_{l,1}^{\top}PD_{l,1})^{-1}(B_{1}^{\top}P+\sum_{l=1}^{r}D_{l,1}^{\top}PC_{lL}+S_{L})=0.$